• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Singh

    Log entry time 02:56:17 on May27,2004

    Entry number 123270

    keyword=septum temp limit & owl shift (2) right septum screenshots

    right septum controls and stripchart @ 0240

    note that during day shift (05/26/04), the "be worried that the septum coil temp is too high" temperature limit was changed to 7.75K from 7.50K after a conversation between Ed and RC Paul. this was done b/c the right septum upper coil temp had been (and still is) stable at about 7.45 ± 0.05K for over 10 hrs yesterday (05/26 owl + 05/25 swing). i've been told that the "software" limit for the septum coil temps are all at 8K, at which point the septum current is ramped down to 0A automatically.


    FIGURE 1

    FIGURE 2